Fermi National Laboratory
 

Operation Summary for August 22, 2011 - August 24, 2011

 

Monday August 22
The day shift began with the Tevatron (TEV) off for a Transfer Hall vacuum leak repair, with Operations monitoring the Recycler stash (300.9E10), with the Antiproton Source (Pbar) stacking (113.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, operators prepared D0 for a 40-minute access at 9:27 AM. FESS personnel reported at 11:02 AM that they had completed cleaning the TEV RF pond water strainer. Operators put the TEV through a dry squeeze at 1:29 PM. Operations established a proton only store at 2:16 PM. Operations established beam to the Meson FTBF T-1015 experiment at 2:44 PM. Operations put the TEV into shot setup at 3:13 PM.

During the evening shift, Operations established store 9017 at 4:23 PM with an average initial luminosity (AIL) of 308.8E30. Operations disabled beam to T-1015 at 8:24 PM.

Tuesday August 23
The midnight shift began with Operations monitoring store 9017 (105.02E30) and the stash (339.3E10), with Pbar stacking (6.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:08 AM, the Pbar beam permit tripped off on a target air flow indication; Operations switched to a NuMI only time line. Operators called a Pbar target person who told them that the target blower had failed; operators called in personnel to repair the blower. Operations terminated store 9017 at 12:45 AM and put the TEV into shot setup. At 1:27 AM, the Booster West Anode Power Supply (WAPS) tripped off; operators responded and reported seeing a flash and hearing a loud bang when the power supply was reset. Operators called in Booster RF personnel to investigate. Operations took the TEV out of shot setup and put it through a dry squeeze at 2:04 AM. Booster RF personnel reported that there was flashover arc that fused the power supply’s door lock. Once they got inside they reported a significant LCW leak and signs of arcing; they began repairs. Operators reported a problem with a Booster power supply (B:MH2) at 3:17 AM; operators tried changing out a regulator card, but that didn’t help. The Pbar person working on the target blower had its motor replaced and the blower back on by 4:41 AM. Operators asked the Booster power supply personnel to look at MH2. Both WAPS and MH2 cam back online by 5:11 AM. At 6:12 AM, WAPS tripped off again; the Booster RF people were still here and investigated; at 6:36 AM, they reported new signs of arcing. Operations put the TEV through a dry squeeze at 6:45 AM. Operations reported at 7 AM that the Neutron Therapy Facility (NTF) began tuning in preparation for treating the first of two patients later this morning. The RF Supply personnel reported that they were changing WAPS’s large resistor at 7:35 AM.

During the day shift, a person working on WAPS reported at 10:34 AM that there was a blown fuse on 13.8kV side of the anode power supply; the High Voltage Coordinator said they would replace the fuse when the rain stopped. Linac personnel reported at 11:10 AM that they had completed their work on LRF5. NTF completed treatment on their last patient for the day at 11:42 AM. At 12:55 PM, LRF5 tripped off for the first of five times in eleven minutes; after the last trip, Linac personnel turned off the station and began repairs. Preaccelerator personnel had the I- Source resistor flushed by 12:56 PM. A Booster person reported at 2 PM that Booster personnel were adding a DI bottle to the WAPS due to its glycol system being contaminated and oxygenated. High Voltage personnel reported at 2:48 PM that they had replaced the fuse for WAPS.

During the evening shift, operators put the TEV through a dry squeeze all shift. WAPS repairs continued all shift.

Wednesday August 24
The midnight shift began with Operations monitoring the stash (329.0E10), with Pbar stacking (10.5E10), and with NuMI and MiniBooNE off due to WAPS glycol skid polishing.

During the midnight shift, at 12:22 AM, operators reported that the Pelletron tripped off; operators reset it and put it in recovery mode. A Pelletron person came in to investigate. Operators reported at 1:22 AM that the resistivity of the WAPS didn’t level off; Booster personnel decided to drain the glycol and fill the system with LCW. Operators called in personnel who had the system drained by 3:28 AM; they began filling it with LCW. Operators prepared MI-31 and the Pelletron for access so that Water Works personnel could begin a back flush of the Pelletron’s solenoid. Operators reported at 5:43 AM that the WAPS was back on; Operations resumed sending beam to Pbar, NuMI, and MiniBooNE. Operators put the TEV through a wet squeeze at 7:07 AM.

The Plans for Wednesday
The plans for today are to stack, stash, and store. However, the Run Coordinator will allow TEV personnel to conduct two 3x3 study stores during the day shift, Water Works personnel will do another flush of the Pelletron’s solenoid, and for MI personnel to conduct an 8 GeV coalescing study. The next HEP store will go in this evening.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/24/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory